calculate the atmospheric pressure if the boiling point of water is 81 degrees Celsius.
1
Expert's answer
2016-09-07T08:51:04-0400
The atmospheric pressure can be calculate with using of Antoine equation: P = 10^(A - B/(C + T)). A, B and C constants for water are 8.07, 1731 and 233. Thus, P = 10^(8.07 - 1731/(233 + 81)) = 363 mmHg.